Marginal and average tax rates Partner A, a single taxpayer, is one of two partners in a small business. As such, she receivePlease answer both A & B of the question! Will show gratitudeX Х More Info Tax Rates-Single Taxpayers-2018 + + $0) + + Range of taxable income $0 to $9,525 9,525 to 38,700 38,700 to 82,5
Answer #1

Answer-a:

Earning
before
taxes ($)
Tax on base income ($) Tax on excess
income ($)
Tax liability ($) After
tax
earning ($)
Avg. tax
rate
13200 upto 9,525 = 953 (13,200 - 9,525)*12% = 441 953 + 441 = 1,394 11806 1,394/13,200*100 = 10.56%
81800 upto 38,700 = 4,454 (81,800 - 38,700)*22% = 9,482 4,454 + 9,482 = 13,936 67864 13,936/81,800*100 = 17.04%
304000 upto 200,000 = 45,690 (304,000 - 200,000)*35% = 36,400 45,690 + 36,400 = 82,090 221910 82,090/304,000*100 = 27%
496000 upto 200,000 = 45,690 (496,000 - 200,000)*35% = 103,600 45,690 + 103,600 = 149,290 346710 149,290/496,000*100 = 30.10%
1200000 upto 500,000 = 150,690 (1,200,000 - 200,000)*37% = 259,000 150,690 + 259,000 = 409,690 790310 409,690/1,200,000*100 = 34.14%
1500000 upto 500,000 = 150,690 (1,500,000 - 200,000)*37% = 370,000 150,690 + 370,000 = 520,690 979310 520,690/1,500,000*100 = 34.71%
1900000 upto 500,000 = 150,690 (1,900,000 - 200,000)*37% = 518,000 150,690 + 518,000 = 668,690 1231310 668,690/1,900,000*100 = 35.19%

Answer-b:

Average tax rate 40.00% 35.00% 30.00% 25.00% 20.00% 15.00% 10.00% 5.00% 0.00% 0 500000 1000000 1500000 2000000 Average tax ra

Comment: As the pre-tax income increases, Avergae tax rate also increase.

Earning before taxes Average tax rate
$                          13,200 10.56%
$                          81,800 17.04%
$                       3,04,000 27.00%
$                       4,96,000 30.10%
$                    12,00,000 34.14%
$                    15,00,000 34.71%
$                    19,00,000 35.19%